Hi!
For homework, I have to solve the equation a(x-h)²+k=0 or a(x-h)^2+k=0 , getting an expression for "x" in terms of the other variables in that equation.
In case this equation looks a little weird, it's the vertex form of a parabola.
I was really stuck on this..but thanks in advance for helping!
*also, please show your work!

Answer by lwsshak3(11628)

solve for x
a(x-h)²+k=0 or a(x-h)^2+k=0
This equation is that of a parabola that opens up or down.
Its standard form: y=a(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex, a=multiplier that affects slope or narrowness of the curve.
..
a(x-h)^2+k=0
a(x-h)^2=-k
(x-h)^2=-k/a
take sqrt of both sides
x-h=±√(-k/a)
x=h±√(-k/a)